OpenStudy (anonymous):

40 divided by X plus 7.5= ? X= -2

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

I am not sure if I understand the question right. Anyway, plug \(x=-2\) in the equation, you get \(\frac{40}{-2+7.5}=\frac{40}{5.5}=\frac{80}{11}\).

OpenStudy (anonymous):

The problem you can get if you write an equation using verbal expressions is that if the equation looks like: \[40/(x) +7.5\] Then by plugging in you would get the answer: \[40/(-2) +7.5 = -12.5\] The way you worded the equation it could go either way...as AnwarA or I stated. Next time try writing the equation exactly how it is given.

OpenStudy (anonymous):

how do we calculate 10 +x divided by 40+ x multiplied by 60%?

OpenStudy (anonymous):

That would be 15% because the x's cancel and you are left with 1/4 which is 25%. Then it's 25% times 60% which is .25 times .6. This results in .15 or 15%.